object a is moving to the right with a velocity of + 5m/s and strikes object B which is initially at rest in an elastic collision. object a and b have equal mass. what is the final velocity of object b


1
Expert's answer
2021-03-14T19:21:00-0400

From the law of conservation of momentum, we have:


"m_au_a=m_av_a+m_bv_b. (1)"

Since collision is elastic, kinetic energy is conserved and we can write:


"\dfrac{1}{2}m_1u_1^2=\dfrac{1}{2}m_1v_1^2+\dfrac{1}{2}m_2v_2^2. (2)"


This formula gives us an additional relationship between velocities. Therefore, with the help of these two formulas we can find the final velocity of object "b" after collision:


"v_b=\dfrac{2m_1u_a}{(m_1+m_2)}=\dfrac{2m_1u_a}{2m_1}=u_a=5\ \dfrac{m}{s}."